HUBBLE SPACE TELESCOPE DAILY REPORT #1522 PERIOD COVERED: 0000Z (UTC) 11/21/95 - 0000Z (UTC) 11/22/95 Daily Status Report as of 326/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/RD 5824 (Occultation of GSC5249 by Saturn) The Faint Object Spectrograph (Red Side) was used to observe = Saturn's atmosphere and its rings while being occulted by GSC5249. = Spectra of the occulted star, of Saturn's atmosphere and of Saturn's rings = will be obtained before and after the occultation. The observations were = completed as planned, and no problems were reported. 1.2 Completed WF/PC-2 5836 (Saturn Ring Plane Crossing Observations in = August and November 1995) The WF/PC-2 was used to make observations of the rings of the = planet Saturn during the ring plan crossing in order to determine their = thickness and distortion. The moons Pan and Atlas will also be observed. = The observations were completed as planned, and no problems were reported. 1.3 Completed WF/PC-2 6250 (WF/PC-2 Cycle 5 Calibration: Internal Monitor) The WF/PC-2 was used to make a series on internal calibration = observations. Internal flats, biases, and K-Spot calibration images were = taken. These observations are designed to monitor the health of the = cameras. The observations were executed as scheduled, and no anomalies = were noted. 1.4 Completed Two Sets of WF/PC-2 6119 (Snapshot Survey for Companions of = Planetary Nebula Nuclei) The WF/PC-2 was used to make snapshot survey observations of = planetary nebula RXJ2117+34 to study its nuclei and to look for possible = companions. The observations were executed as scheduled, and no anomalies = were noted. 1.5 Completed WF/PC-2 5973 (The Host Galaxies of Radio Quite High Redshift = Quasi-stellar Objects) The WF/PC-2 was used to make observations of the host galaxies of = radio quite high red shift quasi-stellar object 1045+060. These data will = be compared to existing data sets for radio-loud, high-redshift QSOs and = ground based data on radio-quiet, low-redshift QSOs. The observations = were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 8 At 326/080939Z, an FGS #3 acquisition failed when the scan step limit was exceeded. HSTAR 5370 describes this event. At 326/102320Z, the guide star acquisition briefly achieved fine lock when FGS #2 experienced a loss of lock. The take data flag was never raised. HSTAR 5371 was written to document this series of events. Scheduled Reacquisitions : 5 Successful : 5 There was a loss of lock at during the second acquisition failure mentioned above. 2.2 FHST Updates: Scheduled : 21 Successful : 21 2.3 Operations Notes: A line outage recorder dump was requested to retrieve 102 seconds = of engineering data lost due to a fatal overflow on the TAC A3 channel. Realtime support was transferred to the SAMS backup facility at = 326/0843Z. This was done to facilitate the examination and testing of the = MOC PRS LAN. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH